T4.025 0 Female undergraduates in randomized groups of 20 took part in a self-esteem s and physical attractiveness. Let x1, x, and x3 be random variables Higher index values mean a more positive infuence on self-esteem study. The study measured an index of self-esteem from the point of view of representing the measure of selif-esteem through x, (competence), 2 (seocial acoeptance), and xy (attractiveness) 20 20 20 18.92 18.64 3.27 3.81 (a) Find a 95% conndence interval for μ1-Pg (Round your answer, to two deomai places.) lower limit upper limit (b) Find a 95% confidence interval for»,-1), (Round your answer, to two donal pace.) lower imit upper limt (E) ,ind a 95% confidence interval for μ2-μ y (Round your answers to two decenal places.) lower Emit upper linma (d) Carernent on the meaning of each org·co hdene intervals found in parts (a). eb), and (e). 서he 95% condence ean you say ab.a N-ve age mnr e competence and attractiveness? between social acceptance and attractiveness nfluence on seif esteem between competence and social acceptance? between
Answer #1

a) 3.272 3.492 μι-,42 = (1892-1864) ± 1.96 20 22.873 20 = 0.28 ± 1.96 = 0.28 ± 2.096 (-1.82, 2.38)

b) 3.272 3.812 in-n-(1892-1798) ±196vAar +- 25.209 20 = 0.94 ± 1.96 = 0.94 ± 2.20 (-1.26, 3.14) c) /3.492 3.812 ,42-μ (18.64-17.98) 1.96 20 20 26.696 20 0.66 ± 1.96 0.66 ± 2.158 (-1.50, 2.82)

Conclusion: here from all three confidence interval we conclude that 2nd and 3rd will be good shows for their corrosponding means diffrrences

at 95 % CI